Miller Company’s contribution format income statement for the most recent month is shown below:

Total Per Unit
Sales (37,000 units) $ 259,000 $ 7.00
Variable expenses 148,000 4.00
Contribution margin 111,000 $ 3.00
Fixed expenses 47,000
Net operating income $ 64,000

Required:

(Consider each case independently):

1. What is the revised net operating income if unit sales increase by 17%?

2. What is the revised net operating income if the selling price decreases by $1.10 per unit and the number of units sold increases by 19%?

3. What is the revised net operating income if the selling price increases by $1.10 per unit, fixed expenses increase by $6,000, and the number of units sold decreases by 7%?

4. What is the revised net operating income if the selling price per unit increases by 20%, variable expenses increase by 30 cents per unit, and the number of units sold decreases by 9%?

Miller Company’s contribution format income statement for the most recent month is shown below:

Total Per Unit
Sales (37,000 units) $ 259,000 $ 7.00
Variable expenses 148,000 4.00
Contribution margin 111,000 $ 3.00
Fixed expenses 47,000
Net operating income $ 64,000

Required:

(Consider each case independently):

1. What is the revised net operating income if unit sales increase by 17%?

2. What is the revised net operating income if the selling price decreases by $1.10 per unit and the number of units sold increases by 19%?

3. What is the revised net operating income if the selling price increases by $1.10 per unit, fixed expenses increase by $6,000, and the number of units sold decreases by 7%?

4. What is the revised net operating income if the selling price per unit increases by 20%, variable expenses increase by 30 cents per unit, and the number of units sold decreases by 9%?

1. NET INCOME

2. NET INCOME

3. NET INCOME

4. NET INCOME
Answer #1

Calculate revised operating income :

1) Revised Net income = (37000*117%*3)-47000 = $82870

2) Revised Net income = (37000*119%*1.9)-47000 = $36657

3) Revised net income = (37000*93%*4.1)-53000 = $88081

4) Revised net income = (37000*91%*4.1)-47000 = $91047